OpenStudy (lena772):

Write an equation of the line through the point (a) parallel to the given line and (b) perpendicular to the given line. point: (-3, 2) line: x+y=7

OpenStudy (lena772):

@Easyaspi314

OpenStudy (anonymous):

For part (a), we want to write the equation of the line through the point (-3,2) and parallel to the line x + y = 7.

OpenStudy (anonymous):

Note, that our line will be parallel to the line x + y = 7, so the slopes MUST be equal.

OpenStudy (lena772):

Okay

OpenStudy (anonymous):

Let's figure out what the slope of the line x+ y = 7 is.

OpenStudy (lena772):

-x?

OpenStudy (anonymous):

write x + y = 7 in the form y = mx+b so we want to solve for y. x + y = 7 subtract x from both sides, we get y = 7 - x So whats the slope of the line? The slope is m, the coefficient of x. So the slope is -1. Agree??????

OpenStudy (lena772):

Yes

OpenStudy (anonymous):

So our line will have a slope of -1.

OpenStudy (lena772):

So the other line must too to be parallel

OpenStudy (lena772):

@Easyaspi314 2=-1(-3)+b?

OpenStudy (anonymous):

So we want to write the equation of the line whose slope is -1 and passing through the point (-3,2). We did write such a line in our last problem, but I will do this one to reinforce the concept. Since our line has slope -1, the equation is y = -1x + b Substitute (-3,2)...that is x = -3 and y = 2 into the equation and solve for b.

OpenStudy (lena772):

b=0?

OpenStudy (anonymous):

So y = -1x + b x = -3, y = 2 so we have 2 = -1(-3) + b 2 = 3 + b b = -1 So the equation of the line will be y = -x - 1.

OpenStudy (lena772):

oh

OpenStudy (anonymous):

To be evry clear, this line y = -x - 1, will pass through (-3,2) and be parallel to the line x+y = 7.

OpenStudy (anonymous):

very*

OpenStudy (anonymous):

Does part (a) make perfect sense???

OpenStudy (lena772):

Yes

OpenStudy (anonymous):

OK..for part (b)...we want to write the equation of the line passing through (-3,2) but perpendicular to the line x + y = 7. Very similar to part(a). You must know that any line perpendicular to the line x+y = 7 will have the negative reciprocal slope. So the negative reciprocal of -2/3 will be +3/2. The line x+y=7 has a slope of -1, as we did before. So the line perpendicular to x+y = 7 will have a slope that is the negative reciprocal of -1. What is the negative reciprocal of -1? It is +1. So our line will have a slope of 1. Now, you can write the equation of the line that has a slope of 1 and passing through

OpenStudy (anonymous):

the point (-3,2).

OpenStudy (anonymous):

You write the equation of the line whose slope is 1 and passing through (-3,2). Then I will see if you are right or wrong.

OpenStudy (anonymous):

ok?

OpenStudy (anonymous):

?

OpenStudy (lena772):

y=x-1?

OpenStudy (anonymous):

Let me see...I didnt do it yet.

OpenStudy (lena772):

ok

OpenStudy (anonymous):

I got y = x+5..let me see what you did

OpenStudy (lena772):

ok

OpenStudy (anonymous):

slope = 1 passing through (-3,2). What did you do to write the equation of the line?

OpenStudy (lena772):

I just changed the slope value to positive i didn't know you had to change the y-intercept

OpenStudy (anonymous):

watch.......

OpenStudy (lena772):

ok

OpenStudy (anonymous):

the general equation of a line is y = mx + b We have a slope of 1, so m = 1 so we have y = 1x + b The question is what is b? so we plug in x = -3 and y = 2 into the above equation...so we get 2 = 1(-3) + b Now we solve for b...we get 22

OpenStudy (anonymous):

froget that 22 at the bottom

OpenStudy (anonymous):

we get 2 = -3 + b b = 5 so the equation of the line will be y = x + 5.

OpenStudy (anonymous):

get it? any questions?

OpenStudy (lena772):

I understand thank you!

OpenStudy (anonymous):

welcome.